CC12:

I think the answer is 24 units What is the area of a rectangle with vertices at ​ (−6, 3) ​, ​ (−3, 6) ​ , (1, 2) , and (−2, −1) ? Enter your answer in the box. Do not round any side lengths.

CC12:

@Eiwoh2 can you help

Eiwoh2:

Sorry, i can't really help with this type of mathematics. I'll get @dude here to help ye. He's a boss. XD

CC12:

Thank you

dude:

D: I’m in mobile so I can’t type all the math but find the distance between 2 points Do you know the distance formula?

CC12:

(-6,3) and (-3,6) d(A,B)=3√2 ≈ 4.2426

CC12:

@dude

dude:

Hmm, I got a different thing, what did you do?

CC12:

\[d(A,B)=\sqrt{(-3-(-6))}^2+(6-3)^2\]

dude:

Ah, all of it goes inside the square root

dude:

\(\sqrt{(6-3)^2+(-3--6)^2}\)

CC12:

9 units

CC12:

sorry it takes me a little longer to figure out the problem i have trouble with math

dude:

Uh, its okay but its not 9 \(\sqrt{(6-3)^2+(-3+6)^2}\\ \sqrt{(3)^2+(3)^2}\\ \sqrt{9+9}\\ \sqrt(18)\approx 4.24\)

dude:

Now find the distance between (-3,6), (1,2)

CC12:

\[\sqrt{1-(-3))^2+(2-6)^2}\] =5.65

CC12:

\[\sqrt{16+16}\] \[\sqrt[4]{2}\]

dude:

Careful, you want to be careful on parenthesis placement \(\sqrt{(2-6)^2+(1--3)^2}\\ \sqrt{(-4)^2+(4)^2}\\ \sqrt{16+16}\\ \sqrt{32}\approx 5.65\)

dude:

Area is length times width, so multiply them

CC12:

23.956